Ahahano408:

HELP PLEASE! IM STRESSING A candy bar box is in the shape of a triangular prism. The volume of the box is 1,200 cubic centimeters. A triangular prism is shown with base of triangle labeled 10 cm, sides of the triangles labeled 13 cm, and length of the box equal to 20 cm. Part A: What is the height of the base? Show your work. (5 points) Part B: What is the approximate amount of cardboard used to make the candy box? Explain how you got your answer. (5 points)

Vocaloid:

part A:  if you split the base into two triangles, you get 2 right triangles with leg 5 and hypotenuse 13. using the pythagorean theorem, letting the height of the triangle be x 5^2 + x^2 = 13^2 solve for x you may also remember that there is a special right triangle involving leg 5 and hypotenuse 13

Vocaloid:

part B) to find the total amount of cardboard, find the surface area of the prism. you need to find the areas of: 1. each triangular base. recall: area of a triangle = (1/2)bh there are 2 triangular bases, so multiply the area of the triangle by 2 2. the area of the left and right rectangular sides, with length 20 and width 13. recall: area of a rectangle = L*W as before, there are 2 sides, so multiply the previous result by 2 3. the area of the bottom rectangle with length 20 and width 10. L*W again. however, this time there is only 1 of these sides. add up the results from steps 1-3